# Evaluating Agents using Social Choice Theory

### 05 Dec 2023 v1

@@ 1 Introduction
To ground our comparison of evaluations methods, we will use the pentathlon depicted in Figure 1 where agents $A$, $B$, and $C$ compete in five different events: archery, basketball, cycling, synchronized swimming, and tennis. Each sport has its own specific evaluation criteria: e.g., fastest time, most points, match outcome. Every sport also awards its top winners with gold, silver, and bronze medals. Who should be considered the winner of the pentathlon? Social choice theory provides a framework for answering this question when we consider the standings within each competitions as votes aggregated according to some voting scheme. Though the voting methods of social choice theory have long been extended beyond the field’s political origins, they have yet to be widely used for the evaluation of machine learning methods.
How do researchers currently evaluate agents? The most commonly used system to compare agents is the Elo rating system (Elo, 1978). Elo was designed originally to rank chess agents. It is still widely-used today in spite of several well-known deficiencies (Balduzzi et al., 2018). For example, it cannot represent non-transitive relationships between agents, as in the classic game of Rock, Paper, Scissors where rock beats scissors, scissors beats paper, and paper beats rock. Because of the cyclic relationship, Elo assigns the same rating to each agent, predicting that one beats the other with probability $\frac{1}{2}$. Elo can also be manipulated by cloning: the addition of multiple copies of similar agents can overturn rankings. A technique such as Nash averaging (Balduzzi et al., 2018) addresses this issue, but raises others. Nash averaging is a game-theoretic approach that interprets evaluation data as a game between agent and task (or agent versus agent), ranking agents by their expected utility against a maximum-entropy Nash equilibrium. We identify three problems with Nash averaging: it requires comparable scores, it aggressively filters out tasks (due to selecting tasks adversarially), and can be highly sensitive to the set of included agents. In contrast, social choice theory provides methods that overcome these problems in a principled manner.
@@
VasE provides several benefits over existing evaluation schemes. VasE does not require score normalization across tasks, only ordinal rankings. This makes its application across incomparable metrics straightforward. Task distribution weights are user-controlled rather than system-controlled. This gives the evaluator the flexibility to decide the relative importance of each task depending on the context. The consistency properties of voting methods are designed to satisfy specific practical axioms, which lead to interpretable comparisons. These axioms also provide guarantees, such as robustness to clones. Finally, compared to Nash averaging, VasE does not filter tasks, and can be more robust to changes in subsets of agents.
@@ 2.0.2 Game-Theoretic Invariant Evaluation
Balduzzi et al. (2018) use the Schur decomposition to separate the problem of agent comparison into cyclic and transitive components, proposing multi-dimensional Elo (mElo) and Nash averaging. mElo can model nontransitive relationships but can be manipulated by adding copies of agents (i.e. it is not clone-proof). In the Agent-vs-Task (AvT) setting, there is a matrix $S$ with $m$ rows (agents) and $n$ columns (tasks) with values corresponding to scores (e.g.normalized average reward) for each agent in each task. In the Agent-vs-Agent (AvA) setting (e.g., Chess), both rows and columns are agents and payoffs are expected utility for each agent. Agents are ranked by their expected values against a maximum-entropy Nash equilibrium strategy of a two-player zero-sum game.
Nash averaging achieves clone-proofness for both agents and tasks due to properties of Nash equilibria. Due to the properties of Nash equilibrium, partial support could be assigned to the task distribution, which we call adversarial task selection. As a result, agents may be assessed over just a few tasks despite a wealth of available data.
Balduzzi et al also argued that an evaluation method should also be invariant if redundant copies of tasks are included. The task invariance property is, however, highly constraining. We show some counter-intuitive effects arising from it in our empirical evaluation section. Furthermore, being invariant to copies of tasks means there is no mechanism to weight the importance of tasks, which can be a useful property. Thus, we do not address frameworks that are invariant in this sense.
@@ 4.1 General Reinforcement Learning Agents
We start with an application to the Arcade Learning Environment (ALE), using the mean performance of eight agents across 54 Atari games from Hessel et al. (2017, Table 5). For VasE, each vote (preference ranking) is determined by decreasing order of score, for a total of 54 complete preference rankings, one for each game. The aggregate rankings (total orders over alternatives) for all nine voting methods are shown in Table 2. Note that all nine methods top-rank Rainbow, which is also the strong Condorcet winner.
@@
Observation 2: Only A3C changes relative ranks across the voting methods. This is evident from the VasE comparison, but is not be immediately obvious from inspection of Hessel et al. (2017, Table 5) nor from the aggregate Elo ratings. This could be due to A3C being the only policy gradient (a fundamentally different category of RL algorithm than DQN and its variants) agent, but suggests further investigation is warranted. Also, A3C has many first-place wins and so is ranked second by plurality and third by STV due to their placing large importance on first-place wins.
Observation 3: The alternate evaluation protocols return identical rankings. ALE is deterministic and so generating unique episodes for comparison requires explicit variation. Hessel et al consider two forms of this: initial periods of forced no-op actions, or taking initial actions taken from human players. Running VasE on the alternative evaluation protocol in Hessel et al. (2017, Table 6) results in identical rankings across all nine voting methods. Thus, the different protocols had insignificant impact on the final assessments.
Observation 4: Nash averaging requires score normalization. Recall that Nash averaging (Balduzzi et al., 2018) finds an equilibrium in the Agent-versus-Task matrix induced by Hessel et al. (2017, Table 5). If we were to run Nash averaging on raw scores, the task player would place full weight on one task (skiing) in its equilibrium because the rewards are large and negative ($<-10000$) and outside the bounds of the other games, across all agents. To avoid this pathology, our Nash Averaging implementation
@@ 4.2 Ranking Large Language Models as Agents
In this section, we apply VasE to the AgentBench benchmark evaluation data (Liu et al., 2023), that assesses the performance of 25 large language models (LLMs), including both APIs and open-source models, as agents across 8 different environments (tasks). The tasks involve operating system and database use, using a knowledge graph to answer questions, playing a digital card game, solving lateral thinking puzzles, house-holding, and browsing and shopping on the web. As in the ALE, VasE builds a preference profile from the raw task scores from Liu et al. (2023, Table 3), without any form of score normalization. In contrast, AgentBench defines an overall score by weighting each agent’s task score by the reciprocal of the task’s mean score.
@@
This scoring system conveys qualitative information about results of each iteration. The top six agents have integer scores: they were Condorcet winners of their respective margin subgames. In other words, the higher-level agents dominate the lower, in terms of performing better on a greater number of tasks. Looking only at the raw data in Liu et al. (2023, Table 3), text-davinci-002 and t-bison-001 seem similar: their overall scores differ only by 0.07 and the scores across tasks are comparable. However, Table 3 above shows these agents perform at different levels. This difference is confirmed by looking at the margin subgame between the bottom three agents: text-davinci-002 is a Condorcet winner of the margin subgame which excludes the top six agents. Though neither Elo nor the aggregate score make it obvious, text-davinci-002 is a stronger performer on these tasks than text-bison-00l and chatglm.
@@ Appendix F Full Results on the Arcade Learning Environment
Table 7: VasE results on Atari with data from Hessel et al. (2017, Table 5) merged with two extra columns (human and random) taken from Badia et al. (2020, Section H.4) .
@@ F.0.3 Analysis of Atari results with $\alpha$-Rank
To try another game-theoretic approach, we also evaluate $\alpha$-rank on the same score matrix. In particular, we use the high ranking-intensity parameter method Omidshafiei et al. (2019, Section 2.5) found in its open reference implementation (Lanctot et al., 2019). In contrast to Nash averaging, $\alpha$-rank finds a full support equilibrium with agent probabilities in $[0.07,0.21]$, task probabilities in $[0.005,0.057]$, and a ranking resembling VasE: rainbow $\succ$ dist-dqn $\succ$ duel $\succ$ prio-ddqn $\succ$ a3c $\succ$ ddqn $\succ$ noisy-dqn $\succ$ dqn. In this case, $\alpha$-rank is considerably less adversarial in the weighting of tasks for ranking. The full weight distributions for agents and tasks found by $\alpha$-rank are given Figures 8 and 7.
@@ Appendix G Full Results on Adaptive Agents
  • 119: An agent that used "skip memory" as described in Team et al. (2023, Section 3.7), combining a GRU and Transformer XL to increase the effective context length.
@@ 10 Cohere xlarge v20220609 52B 514
="ltx_page_logo">Generated on Tue Dec 5 20:39:46 2023 by

### 07 Dec 2023 v2

@@ 1 Introduction
To ground our comparison of evaluations methods, we will use the pentathlon depicted in Figure 1 where agents $A$, $B$, and $C$ compete in five different events: archery, basketball, cycling, synchronized swimming, and tennis. Each sport has its own specific evaluation criteria: e.g., fastest time, most points, match outcome. Every sport also awards its top winners with gold, silver, and bronze medals. Who should be considered the winner of the pentathlon? Social choice theory provides a framework for answering this question when we consider the standings within each competitions as votes aggregated according to some voting scheme. Though the voting methods of social choice theory have long been extended beyond the field’s political origins, they have yet to be widely used for the evaluation of machine learning agents.
How do researchers currently evaluate agents? The most commonly used system to compare agents is the Elo rating system (Elo, 1978). Elo was designed originally to rank chess agents. It is still widely-used today in spite of several well-known deficiencies (Balduzzi et al., 2018). For example, it cannot represent non-transitive relationships between agents, as in the classic game of Rock, Paper, Scissors where rock beats scissors, scissors beats paper, and paper beats rock. Because of the cyclic relationship, Elo assigns the same rating to each agent, predicting that one beats the other with probability $\frac{1}{2}$. Elo can also be manipulated by cloning: the addition of multiple copies of similar agents can overturn rankings. A technique such as Nash averaging (Balduzzi et al., 2018) addresses the cloning issue, but raises others. Nash averaging is a game-theoretic approach that interprets evaluation data as a game between agent and task (or agent versus agent), ranking agents by their expected utility against a maximum-entropy Nash equilibrium. We identify three problems with Nash averaging: it requires scores be directly comparable, it aggressively filters out tasks (due to selecting tasks adversarially), and can be highly sensitive to the set of included agents. In contrast, social choice theory provides methods that overcome these problems in a principled manner.
@@
VasE provides several benefits over existing evaluation schemes. VasE does not require score normalization across tasks, only ordinal rankings. This makes it straightforward to apply, even across incomparable metrics. Task distribution weights are user-controlled rather than system-controlled. This gives the evaluator the flexibility to decide the relative importance of each task depending on the context. The aggregate outcome is interpretable, as the consistency properties of voting methods are designed to satisfy specific practical axioms. These axioms also provide guarantees, such as robustness to clones. Finally, compared to Nash averaging, VasE does not filter tasks, and can be more robust to changes in subsets of agents.
@@ 2.0.2 Game-Theoretic Invariant Evaluation
Balduzzi et al. (2018) use the Schur decomposition to separate the problem of agent comparison into cyclic and transitive components, proposing multi-dimensional Elo (mElo) and Nash averaging. In the Agent-vs-Task (AvT) setting, they use a matrix $S$ with $m$ rows (agents) and $n$ columns (tasks) with values corresponding to scores (e.g.normalized average reward) for each agent in each task. In the Agent-vs-Agent (AvA) setting (e.g., Chess), both rows and columns are agents and payoffs are expected utility for each agent. Agents are ranked by their expected values against a maximum-entropy Nash equilibrium strategy of a two-player zero-sum game. mElo can model nontransitive relationships but can be manipulated by adding copies of agents (i.e. it is not clone-proof).
Nash averaging achieves clone-proofness for both agents and tasks due to properties of Nash equilibria. However, these properties allow partial support to be assigned to the task distribution, which we call adversarial task selection. As a result, agents may be assessed over just a few tasks despite a wealth of available data.
Balduzzi et al. (2018) argue that an evaluation method should also be invariant to redundant copies of tasks. This task invariance property is, however, highly constraining. We show some counter-intuitive effects arising from it in our empirical evaluation section. Furthermore, being invariant to copies of tasks means there is no mechanism to weight the importance of tasks, which is often be a useful feature. Thus, we do not address frameworks that are invariant in this sense.
@@ 4.1 General Reinforcement Learning Agents
We start with an application to the Arcade Learning Environment (ALE), using the mean performance of eight agents across 54 Atari games from (Hessel et al., 2017, Table 5). For VasE, each vote (preference ranking) is determined by decreasing order of score, for a total of 54 complete preference rankings, one for each game. The aggregate rankings (total orders over alternatives) for all nine voting methods are shown in Table 2. Note that all nine methods top-rank Rainbow, which is also the strong Condorcet winner.
@@
Observation 2: Only A3C changes relative ranks across the voting methods. This is evident from the VasE comparison, but is not be immediately obvious from inspection of (Hessel et al., 2017, Table 5) nor from the aggregate Elo ratings. This could be due to A3C being the only policy gradient (a fundamentally different category of RL algorithm than DQN and its variants) agent, but suggests further investigation is warranted. Also, A3C has many first-place wins and so is ranked second by plurality and third by STV due to their placing large importance on first-place wins.
Observation 3: The alternate evaluation protocols return identical rankings. ALE is deterministic and so generating unique episodes for comparison requires explicit variation. Hessel et al consider two forms of this: initial periods of forced no-op actions, or taking initial actions taken from human players. Running VasE on the alternative evaluation protocol in (Hessel et al., 2017, Table 6) results in identical rankings across all nine voting methods. Thus, the different protocols had insignificant impact on the final assessments.
Observation 4: Nash averaging requires score normalization. Recall that Nash averaging (Balduzzi et al., 2018) finds an equilibrium in the Agent-versus-Task matrix induced by (Hessel et al., 2017, Table 5). If we were to run Nash averaging on raw scores, the task player would place full weight on one task (skiing) in its equilibrium because the rewards are large and negative ($<-10000$) and outside the bounds of the other games, across all agents. To avoid this pathology, our Nash Averaging implementation
@@ 4.2 Ranking Large Language Models as Agents
In this section, we apply VasE to the AgentBench benchmark evaluation data (Liu et al., 2023), that assesses the performance of 25 large language models (LLMs), including both APIs and open-source models, as agents across 8 different environments (tasks). The tasks involve operating system and database use, using a knowledge graph to answer questions, playing a digital card game, solving lateral thinking puzzles, house-holding, and browsing and shopping on the web. As in the ALE, VasE builds a preference profile from the raw task scores from (Liu et al., 2023, Table 3), without any form of score normalization. In contrast, AgentBench defines an overall score by weighting each agent’s task score by the reciprocal of the task’s mean score.
@@
This scoring system conveys qualitative information about results of each iteration. The top six agents have integer scores: they were Condorcet winners of their respective margin subgames. In other words, the higher-level agents dominate the lower, in terms of performing better on a greater number of tasks. Looking only at the raw data in (Liu et al., 2023, Table 3), text-davinci-002 and t-bison-001 seem similar: their overall scores differ only by 0.07 and the scores across tasks are comparable. However, Table 3 above shows these agents perform at different levels. This difference is confirmed by looking at the margin subgame between the bottom three agents: text-davinci-002 is a Condorcet winner of the margin subgame which excludes the top six agents. Though neither Elo nor the aggregate score make it obvious, text-davinci-002 is a stronger performer on these tasks than text-bison-00l and chatglm.
@@ Appendix F Full Results on the Arcade Learning Environment
Table 7: VasE results on Atari with data from (Hessel et al., 2017, Table 5) merged with two extra columns (human and random) taken from (Badia et al., 2020, Section H.4) .
@@ F.0.3 Analysis of Atari results with $\alpha$-Rank
To try another game-theoretic approach, we also evaluate $\alpha$-rank on the same score matrix. In particular, we use the high ranking-intensity parameter method (Omidshafiei et al., 2019, Section 2.5) found in its open reference implementation (Lanctot et al., 2019). In contrast to Nash averaging, $\alpha$-rank finds a full support equilibrium with agent probabilities in $[0.07,0.21]$, task probabilities in $[0.005,0.057]$, and a ranking resembling VasE: rainbow $\succ$ dist-dqn $\succ$ duel $\succ$ prio-ddqn $\succ$ a3c $\succ$ ddqn $\succ$ noisy-dqn $\succ$ dqn. In this case, $\alpha$-rank is considerably less adversarial in the weighting of tasks for ranking. The full weight distributions for agents and tasks found by $\alpha$-rank are given Figures 8 and 7.
@@ Appendix G Full Results on Adaptive Agents
  • 119: An agent that used "skip memory" as described in (Team et al., 2023, Section 3.7), combining a GRU and Transformer XL to increase the effective context length.
@@ 10 Cohere xlarge v20220609 52B 514
="ltx_page_logo">Generated on Thu Dec 7 02:17:16 2023 by